OpenStudy (anonymous):

A wooden fence 6 ft. high and 300 ft. long is to be painted on one side. How many gallons of paint are needed if one gallon covers 400 sq. ft.?

OpenStudy (anonymous):

area of fence = 6 * 300 = 1800 ft^2

OpenStudy (anonymous):

No of gallons required = 1800/400 = 4.5

OpenStudy (anonymous):

First, you should calculate the square footage of the fence.\[6 ft * 300 ft = 1800 ft ^{2}\] Then to find the number of gallons of paint needed, divide the area of the fence by the area covered with each gallon of paint: \[1800 ft ^{2} \div 400 ft ^{2}= 4.5 \] So, 4.5 gallons of paint, or rounding up, 5.